OpenStudy (kaylak12345):

simplify 2 square root of 5 3 square root of 7^ 2

OpenStudy (kaylak12345):

simplify 4 square root of 6 over square root of 30 by rationalizing the denominator

OpenStudy (studygurl14):

Only one question at a time. Is the first one \(\Large (2\sqrt{5})(3\sqrt{7^2}\)

OpenStudy (studygurl14):

i meant to put another parenthesis at the end.

OpenStudy (kaylak12345):

no (\[(2\sqrt{5}+3\sqrt{7})^2\]

OpenStudy (kaylak12345):

is the answer 2*2*squrt5*3*squrt7=2*2*3*squrt5*squrt7

OpenStudy (studygurl14):

Oh, okay. So you have \(\Large (2\sqrt{5}+3\sqrt{7})(2\sqrt{5}+3\sqrt{7})\)

OpenStudy (kaylak12345):

okay

OpenStudy (studygurl14):

If \(\Large (a+b)(a+b)=a^2+2ab+b^2\) and a=\(\large 2\sqrt{3}\) and b=\(\large 3\sqrt{7}\), what does this expression equal?

OpenStudy (kaylak12345):

\[4\sqrt{3}+6\sqrt{7}=2\sqrt{3}^2+4\sqrt{3}+6\sqrt{7}+3\sqrt{7}^2\]

OpenStudy (kaylak12345):

something like that idk

OpenStudy (studygurl14):

Not quite. (sorry i made a mistake earlier, it was \(\sqrt{5}\), not \(\sqrt{3}\))

OpenStudy (kaylak12345):

im so confused offically

OpenStudy (studygurl14):

hold on a sec

OpenStudy (studygurl14):

\(\Large a^2+2ab+b^2\) \(\Large a=2\sqrt{5}\), \(\Large b=3\sqrt{7}\) \(\Large (2\sqrt{5})^2+2(2\sqrt{5})(3\sqrt{7})+(3\sqrt{7})^2\)

OpenStudy (kaylak12345):

so is that the answer of what i have to solve

OpenStudy (studygurl14):

\(\large \rightarrow 2^2(\sqrt{5})^2+2(2\sqrt{5})(3\sqrt{7})+3^2(\sqrt{7})^2\rightarrow?\)

OpenStudy (kaylak12345):

so thta is what i am solving

OpenStudy (studygurl14):

yeah. simplify

OpenStudy (kaylak12345):

first one is 2squrt 5^4

OpenStudy (studygurl14):

No. First off, what is 2^2?

OpenStudy (kaylak12345):

4

OpenStudy (studygurl14):

correct. Now what is (\sqrt{5})^2\)

OpenStudy (studygurl14):

\((\sqrt{5}^2\)

OpenStudy (studygurl14):

darnnit, lol. i keep messing it up

OpenStudy (kaylak12345):

my calculator says 5

OpenStudy (kaylak12345):

is it 5

OpenStudy (kaylak12345):

is it 5

OpenStudy (kaylak12345):

hello

OpenStudy (studygurl14):

yep 5

OpenStudy (kaylak12345):

ok

OpenStudy (studygurl14):

Okay, so you get... \(\Large 5 + 2(2\sqrt{5})(3\sqrt{7})+3^2(\sqrt{7})^2\)

OpenStudy (studygurl14):

What's 3^2? What's (sqrt{7})^2 ?

OpenStudy (kaylak12345):

9 and 7

OpenStudy (kaylak12345):

@Tlynn514

OpenStudy (kaylak12345):

can u help me finidh

OpenStudy (kaylak12345):

@robtobey

OpenStudy (kaylak12345):

or what is the answer so i can check my work

OpenStudy (anonymous):

I am so confused...I read the above equations that other people put up and I keep doing the problem wrong.

OpenStudy (anonymous):

2×2×5√×3×7√=2×2×3×5√×7√ =12×5√×7√

OpenStudy (kaylak12345):

type it please

OpenStudy (anonymous):

refresh the page and it will come across clearly

OpenStudy (anonymous):

Did it work?

OpenStudy (kaylak12345):

so is the answer 83+12 square root of 35

OpenStudy (anonymous):

YESH!

OpenStudy (kaylak12345):

yay

OpenStudy (anonymous):

:)